Solve in which the triangle a=(√3+1),b=(√3-1)and c=60°
 

Grade:12

1 Answers

Arun
5 years ago

Let C= 60o ⇒ B + A = 120o.

Also tan (A – B)/ 2 = (a – b)/(a + b) cot C/2

= (√3 + 1 - √3 + 1) / (√3 + 1 + √3 – 1) cot 30o 

= 1

⇒ (A – B)/2 = 45o 

⇒ A – B = 90o

⇒ A = 105o, B = 15o

Also c = a sin  / sin A = (√3 + 1)/ sin 60° cos 45° + cos 60° sin 45°

= √2(√3 + 1). √3/(√3 + 1) = √6.